OpenStudy (anonymous):

A million years ago, an alien species built a vertical tower on a horizontal plane. When they returned they discovered that the ground had tilted so that measurements of 3 points on the ground gave coordinates of (0, 0, 0), (2, 1, 0), and (0, 3, 1). By what angle does the tower now deviate from the vertical?

OpenStudy (amistre64):

 seems like an accurate depiction, define 2 vectors and take the cross product to determine the "normal to the plane" the angle this makes with the 0,0,1 vector tells how far it has leaned

OpenStudy (amistre64):

if they coords they give represent abc, then bxc should give us a normal to compare with

OpenStudy (amistre64):

since the z axis of 0,0,1 has a length of 1 ... \[cos\alpha=\frac{\vec n\cdot (0,0,1)}{|n|}\] \[ cos\alpha=\frac{\vec n_{Z}}{|n|}\]

OpenStudy (amistre64):

can you work that out?

OpenStudy (anonymous):

I got (1,2,6) as the normal, do I get the dot product of n and (0,0,1) and divide it by the length of the normal?

OpenStudy (amistre64):

yes, which you will soon find out is just the z component of the unit normal

OpenStudy (anonymous):

1/7

OpenStudy (amistre64):

6/sqrt(1+4+36) 6/sqrt(41)

OpenStudy (anonymous):

sorry I forgot to square.. do I divide 6/sqrt(41) by length of n?

OpenStudy (amistre64):

no, the length of n is sqrt41 the z component of the normal ... if we dot (x,y,z) with (0,0,1) we get 0+0+z = z the z component is 6, the length is sqrt41

OpenStudy (anonymous):

do I use the cosine rule?

OpenStudy (amistre64):

id hope so ... since cos(angle) = 6/sqrt(41), arc it

OpenStudy (anonymous):

0.3567333885

OpenStudy (amistre64):

you set to radians or degrees?

OpenStudy (anonymous):

radians, I'm looking for the answer in radians

OpenStudy (anonymous):

it is correct, Thank you vey much for your help

OpenStudy (amistre64):

making me have to reset the mode on my ti83 ... i should have you flogged!! :)